# Structured Symmetric Tensors

### v1

@@ Abstract
In this paper, we study structured symmetric tensors. We introduce several new classes of structured symmetric tensors: completely decomposable (CD) tensors, strictly sum of squares (SSOS) tensors and SOS tensors. CD tensors have applications in data analysis and signal processing. Complete Hankel tensors are CD tensors. SSOS tensors are defined as SOS tensors with a positive definite Gram matrix, ensuring structural stability under perturbations. The SOS cone is defined as the dual cone of the SOS tensor cone, with characterizations via moment matrices and polynomial nonnegativity. We study the relations among completely positive (CP) cones, CD cones, sum of squares (SOS) cones, positive semidefinite (PSD) cones and copositive (COP) cones. We identify the interiors of PSD, SOS, CP, COP and CD cones for even-order tensors. These characterizations are crucial for interior-point methods and stability analysis in polynomial and tensor optimization. We generalize the classical Schur product theorem to CD and CP tensors, including the case of strongly completely decomposable (SCD) and strongly completely positive (SCP) tensors. We identify equivalence between strictly CD (SCD) and positive definite (PD) for CD tensors. This result establishes a key link between the spanning property of decomposition vectors and positive definiteness.
Key words. Completely decomposable tensors, positive semidefinite tensors, sum of squares tensors, completely positive tensors, copositive tensors, interiors, dual cones.
@@ 1 Introduction
Symmetric tensors have important applications in multi-variable calculus, multi-variate polynomial optimization, signal processing, magnetic resonance imaging and spectral hypergraph theory [5, Section 1.1]. Several classes of structured tensors have already been studied extensively in the literature. These include positive semidefinite (PSD) tensors, sum of squares (SOS) tensors, completely positive (CP) tensors and copositive (COP) tensors. Actually, the above four classes of structured symmetric tensors form closed and convex cones in the corresponding symmetric tensor spaces. In particular, the CP and COP cones are dual cones to each other [5, 6].
@@
In Section 5, we study CD tensors more. We identify equivalence between strictly CD (SCD) and positive definite (PD) for CD tensors. This result establishes a key link between the spanning property of decomposition vectors and positive definiteness. We then identify completely Hankel tensors are CD tensors. Since Hankel tensors have wide applications in signal processing and data analysis [5], this justifies the importance of CD tensors.
@@ 2 Important Structured Symmetric Tensor Classes
$\mathcal{A}\mathbf{x}^{m}\equiv\sum_{j=1}^{r}(C_{j}\mathbf{x}^{k})^{2},$
then we say that $\mathcal{A}$ is an sum-of-squares (SOS) tensor. The smallest $r$ is called the SOS rank of $\mathcal{A}$. Clearly, if $\mathcal{A}$ is an SOS tensor, then $\mathcal{A}$ is PSD, but not vice versa in general. All the SOS tensors in $S_{m,n}$ also form a closed and convex cone. Denote it by $SOS_{m,n}$. Then $SOS_{m,n}\subset PSD_{m,n}$. By Hilbert [1], $SOS_{m,n}=PSD_{m,n}$ only in the following three cases: (i) $m=2$, (ii) $n=2$, (iii) $m=4$ and $n=3$.
@@
$\mathcal{A}=\sum_{j=1}^{p}\mathbf{u}_{j}^{m},$
then $\mathcal{A}$ is called a completely decomposable (CD) tensor. The smallest value of $p$ is called the CD rank of $\mathcal{A}$. If furthermore, $\mathbf{u}_{j},j\in[p]$ spans $\mathds{R}^{n}$, then we say that $\mathcal{A}$ is a strongly completely decomposable (SCD) tensor. Denote the set of all the CD tensors in $S_{m,n}$ by $CD_{m,n}$, and the set of all the SCD tensors in $S_{m,n}$ by $SCD_{m,n}$. Then $CD_{m,n}$ is also a closed and convex cone. If $m$ is odd, then $CD_{m,n}=S_{m,n}$. On the other hand, if $m=2k$, then a CD tensor is an SOS tensor, i.e., $CD_{m,n}\subset SOS_{m,n}$, as in this case, for $\mathbf{x}\in\mathds{R}^{n}$, we have
$\mathcal{A}\mathbf{x}^{m}=\sum_{j=1}^{p}(\mathbf{u}_{j}^{k}\bullet\mathbf{x}^{k})^{2}.$
We have $CD_{m,n}\subsetneq SOS_{m,n}$. For example, for $m=4$ and $n=2$, $\mathcal{A}\mathbf{x}^{4}=(x_{1}^{2}+x_{2}^{2})^{2}$ is SOS, but not CD. When $m$ is even, by [3], $CD_{m,n}=PSD_{m,n}^{*}$ and $PSD_{m,n}=CD_{m,n}^{*}$.
@@
(iii) The inclusion $CP_{m,n}\subset CD_{m,n}$ is strict because CP tensors require nonnegative vectors while CD tensors allow any real vectors. The inclusion $CD_{m,n}\subset SOS_{m,n}$ follows from the representation $\mathcal{A}\mathbf{x}^{m}=\sum_{j=1}^{p}(\mathbf{u}_{j}^{k}\bullet\mathbf{x}^{k})^{2}$ when $m=2k$, and is strict as shown by the example $\mathcal{A}\mathbf{x}^{4}=(x_{1}^{2}+x_{2}^{2})^{2}$ for $m=4,n=2$. The inclusion $SOS_{m,n}\subset PSD_{m,n}$ is by definition, and $PSD_{m,n}\subset COP_{m,n}$ since nonnegativity on all $\mathbf{x}\in\mathds{R}^{n}$ implies nonnegativity on $\mathds{R}^{n}_{+}$. Both inclusions are strict in general.
(iv) The equality cases for SOS and PSD tensors are classical results from Hilbert [1]. For $m=2$, CD tensors coincide with SOS tensors because any PSD matrix can be written as a sum of squares of linear forms. For $m\geq 4$, the strict inclusion $CD_{m,n}\subsetneq SOS_{m,n}$ is demonstrated by the biquadratic form $(x_{1}^{2}+x_{2}^{2})^{2}$ which is SOS but not CD. ∎
@@ 3 Interiors of PSD, SOS, CP, COP and CD Cones
  • $CD_{m,n}=S_{m,n}$, so its interior is $S_{m,n}$ itself, while $SCD_{m,n}$ is a proper subset.
@@
  • $SOS_{m,n}$ is not defined for odd $m$ since the SOS representation requires even order.
@@ 4.1 Schur Product Theorem for CD Tensors
Since $k$ is arbitrary, the set $\mathbf{u}_{i}\circ\mathbf{v}_{j}:i\in[p],j\in[q]$} generates all coordinate vectors. Hence, $\mathcal{A}\circ\mathcal{B}$ is also SCD.
@@ 4.2 Schur Product Theorem for CP Tensors
(i) and (iii) are Proposition 6.5 of [5]. The proof of (ii) is similar to the proof of Theorem 4.1 (ii).
@@ 6.1 Answers to Fundamental Questions
    No. The SOS cone is not self-dual in general. For example, when $m=4$ and $n=2$, we have $SOS_{4,2}=PSD_{4,2}$ and $SOS_{4,2}^{*}=CD_{4,2}$. The tensor $\mathcal{A}$ defined by $\mathcal{A}\mathbf{x}^{4}=(x_{1}^{2}+x_{2}^{2})^{2}$ is SOS but not completely decomposable (CD), hence not SOS.
@@ 7 Final Remarks
Acknowledgment This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), the R&D project of Pazhou Lab (Huangpu) (Grant no. 2023K0603), the Fundamental Research Funds for the Central Universities (Grant No. YWF-22-T-204), Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002).

### v2

@@ Abstract
In this paper, we study structured symmetric tensors. We introduce several new classes of structured symmetric tensors: completely decomposable (CD) tensors, strictly sum of squares (SSOS) tensors and SOS tensors. CD tensors have applications in data analysis and signal processing. Complete Hankel tensors are CD tensors. SSOS tensors are defined as SOS tensors with a positive definite Gram matrix, ensuring structural stability under perturbations. The SOS cone is defined as the dual cone of the SOS tensor cone, with characterizations via moment matrices and polynomial nonnegativity. We study the relations among completely positive (CP) cones, CD cones, sum of squares (SOS) cones, positive semidefinite (PSD) cones and copositive (COP) cones. We identify the interiors of PSD, SOS, CP, COP and CD cones for even-order tensors. These characterizations are crucial for interior-point methods and stability analysis in polynomial and tensor optimization. We generalize the classical Schur product theorem to CD and CP tensors, including the case of strongly completely decomposable (SCD) and strongly completely positive (SCP) tensors. We identify equivalence between strictly CD (SCD) and positive definite (PD) for CD tensors. Furthermore, we give an example of a PSD but not SOS Hankel tensor. This answers an open question raised in the literature.
Key words. Completely decomposable tensors, positive semidefinite tensors, sum of squares tensors, Hankel tensors, interiors, dual cones.
@@ 1 Introduction
Symmetric tensors have important applications in multi-variable calculus, multi-variate polynomial optimization, signal processing, magnetic resonance imaging and spectral hypergraph theory [8, Section 1.1]. Several classes of structured tensors have already been studied extensively in the literature. These include positive semidefinite (PSD) tensors, sum of squares (SOS) tensors, completely positive (CP) tensors and copositive (COP) tensors [8, 10]. Actually, the above four classes of structured symmetric tensors form closed and convex cones in the corresponding symmetric tensor spaces. In particular, the CP and COP cones are dual cones to each other [8, 9].
@@
In Section 5, we study CD tensors more. We identify equivalence between strictly CD (SCD) and positive definite (PD) for CD tensors. This result establishes a key link between the spanning property of decomposition vectors and positive definiteness. We then identify completely Hankel tensors are CD tensors. Since Hankel tensors have wide applications in signal processing and data analysis [8], this justifies the importance of CD tensors. We further show that the inheritance property of Hankel tensors also holds for the CD property. Furthermore, we give an example of a PSD but not SOS Hankel tensor. This answers an open question raised in Subsection 5.7.10 of [8].
@@ 2 Important Structured Symmetric Tensor Classes
$\mathcal{A}\mathbf{x}^{m}\equiv\sum_{j=1}^{r}\left({\mathcal{C}_{j}\mathbf{x}^{k}}\right)^{2},$
then we say that $\mathcal{A}$ is a sum-of-squares (SOS) tensor. The smallest $r$ is called the SOS rank of $\mathcal{A}$. Clearly, if $\mathcal{A}$ is an SOS tensor, then $\mathcal{A}$ is PSD, but not vice versa in general. All the SOS tensors in $S_{m,n}$ also form a closed and convex cone. Denote it by $SOS_{m,n}$. Then $SOS_{m,n}\subset PSD_{m,n}$. By Hilbert [3], $SOS_{m,n}=PSD_{m,n}$ only in the following three cases: (i) $m=2$, (ii) $n=2$, (iii) $m=4$ and $n=3$.
@@
$\mathcal{A}=\sum_{j=1}^{p}\mathbf{u}_{j}^{m},$ (1)
then $\mathcal{A}$ is called a completely decomposable (CD) tensor. The smallest value of $p$ is called the CD rank of $\mathcal{A}$. If furthermore, $\mathbf{u}_{j},j\in[p]$ spans $\mathds{R}^{n}$, then we say that $\mathcal{A}$ is a strongly completely decomposable (SCD) tensor. Denote the set of all the CD tensors in $S_{m,n}$ by $CD_{m,n}$, and the set of all the SCD tensors in $S_{m,n}$ by $SCD_{m,n}$. Then $CD_{m,n}$ is also a closed and convex cone. If $m$ is odd, then (1) is equivalent to the symmetric outer product decomposition. Thus, it follows from [2] that $CD_{m,n}=S_{m,n}$ and $p\leq\binom{n+m-1}{m}$. On the other hand, if $m=2k$, then a CD tensor is an SOS tensor, i.e., $CD_{m,n}\subset SOS_{m,n}$, as in this case, for $\mathbf{x}\in\mathds{R}^{n}$, we have
$\mathcal{A}\mathbf{x}^{m}=\sum_{j=1}^{p}\left({\mathbf{u}_{j}^{k}\bullet\mathbf{x}^{\otimes k}}\right)^{2}.$
We have $CD_{m,n}\subsetneq SOS_{m,n}$. For example, for $m=4$ and $n=2$, a Hankel tensor in Subsection 5.4 is SOS, but not CD. When $m$ is even, by [5], $CD_{m,n}=PSD_{m,n}^{*}$ and $PSD_{m,n}=CD_{m,n}^{*}$. In 2017, Nie [6] introduced a generating polynomial method for computing symmetric tensor decompositions, which is applicable to CD decompositions in the complex field. The real field, in contrast, still lacks an effective numerical procedure to verify whether a tensor is a CD tensor and to subsequently compute its decomposition.
@@
(iii) The inclusion $CP_{m,n}\subset CD_{m,n}$ is strict because CP tensors require nonnegative vectors while CD tensors allow any real vectors. The inclusion $CD_{m,n}\subset SOS_{m,n}$ follows from the representation $\mathcal{A}\mathbf{x}^{m}=\sum_{j=1}^{p}(\mathbf{u}_{j}^{k}\bullet\mathbf{x}^{k})^{2}$ when $m=2k$, and is strict as shown by a Hankel tensor example in Subsection 5.4 for $m=4,n=2$. The inclusion $SOS_{m,n}\subset PSD_{m,n}$ is by definition, and $PSD_{m,n}\subset COP_{m,n}$ since nonnegativity on all $\mathbf{x}\in\mathds{R}^{n}$ implies nonnegativity on $\mathds{R}^{n}_{+}$. Both inclusions are strict in general.
(iv) The equality cases for SOS and PSD tensors are classical results from Hilbert [3]. For $m=2$, CD tensors coincide with SOS tensors because any PSD matrix can be written as a sum of squares of linear forms. For $m\geq 4$, the strict inclusion $CD_{m,n}\subsetneq SOS_{m,n}$ is demonstrated by the Hankel tensor example in Subsection 5.4, which is SOS but not CD. ∎
@@ 3 Interiors of PSD, SOS, CP, COP and CD Cones
  • $CD_{m,n}=S_{m,n}$, so its interior is the empty set.
@@
  • $SOS_{m,n}$ is not defined for odd $m$ since the SOS representation requires even order
@@ 4.1 Schur Product Theorem for CD Tensors
Since $k$ is arbitrary, the set $\mathbf{u}_{i}\circ\mathbf{v}_{j}:i\in[p],j\in[q]$} generates all coordinate vectors. Hence, $\mathcal{A}\circ\mathcal{B}$ is also SCD. ∎
@@ 4.2 Schur Product Theorem for CP Tensors
(i) and (iii) are Proposition 6.5 of [8]. The proof of (ii) is similar to the proof of Theorem 4.1 (ii). ∎
@@ 5.2 Complete Hankel Tensors
5.3 The Inheritance Property of Hankel Tensors
Hankel tensors has an inheritance property. Let $\mathcal{A}=(a_{i_{1}\cdots i_{m}})\in S_{m,n}$ be a Hankal tensor with a generating vector ${\bf h}\in\mathds{R}^{(n-1)m+1}$. Let n1=(p1)qn-1=(p-1)q, where $q\geq 2$ and $p\geq 3$. Then $\bf h$} is also a generating vector of a Hankel tensor $\mathcal{B}=(b_{j_{1}\cdots j_{qm}}\in S_{qm,p}$, with
$b_{j_{1}\cdots j_{qm}}=h_{i_{1}+\cdots+i_{qm}-n+1},$ (6)
for $j_{1},\cdots,j_{qm}\in[p]$. It was shown in [8, Theorem 5.53] that $\mathcal{B}$ is PSD, or PD or SOS if $\mathcal{A}$ is PSD, or PD or SOS, respectively.
Theorem 5.2.
Under the above setting, if $\mathcal{A}$ is CD, then $\mathcal{B}$ is CD.
Proof.
Since $\mathcal{A}$ is a CD tensor and a Hankel tensor, it is a complete Hankel tensor. Therefore, $\mathcal{A}$ admits a Vandermonde decomposition with nonnegative coefficients:
$\mathcal{A}=\sum_{j=1}^{r}\alpha_{j}(\mathbf{u}_{j})^{m},$
where $\alpha_{j}\geq 0$ and $\mathbf{u}_{j}=(1,u_{j},u_{j}^{2},\ldots,u_{j}^{n-1})^{\top}$ are Vandermonde vectors in $\mathds{R}^{n}$.
Now consider the Hankel tensor $\mathcal{B}\in S_{qm,p}$ defined by the same generating vector $\bf h$}. For each Vandermonde vector $\mathbf{u}_{j}\in\mathds{R}^{n}$, define the corresponding Vandermonde vector $\mathbf{w}_{j}\in\mathds{R}^{p}$ by:
$\mathbf{w}_{j}=(1,u_{j}^{q},u_{j}^{2q},\ldots,u_{j}^{(p-1)q})^{\top}.$
Note that since n1=(p1)qn-1=(p-1)q, the exponents in $\mathbf{w}_{j}$ exactly match the required range.
We claim that $\mathcal{B}$ has the Vandermonde decomposition:
$\mathcal{B}=\sum_{j=1}^{r}\alpha_{j}(\mathbf{w}_{j})^{qm}.$
To verify this, consider the generating vector $\bf h$}. From the Vandermonde decomposition of $\mathcal{A}$, we have:
$h_{i}=\sum_{j=1}^{r}\alpha_{j}u_{j}^{i-1},\qquad i=1,\ldots,(n-1)m+1.$
For $\mathcal{B}$, the entry $b_{j_{1}\cdots j_{qm}}$ is determined by:
$b_{j_{1}\cdots j_{qm}}=h_{j_{1}+\cdots+j_{qm}-p+1}.$
On the other hand, the corresponding entry in $\sum_{j=1}^{r}\alpha_{j}(\mathbf{w}_{j})^{qm}$ is:
$\sum_{j=1}^{r}\alpha_{j}\left(u_{j}^{q(j_{1}-1)}\cdots u_{j}^{q(j_{qm}-1)}\right)=\sum_{j=1}^{r}\alpha_{j}u_{j}^{q(j_{1}+\cdots+j_{qm}-qm)}.$
Since $j_{1}+\cdots+j_{qm}-p+1=q(j_{1}+\cdots+j_{qm}-qm)$ when n1=(p1)qn-1=(p-1)q, the two expressions are equal. Therefore, $\mathcal{B}$ admits a Vandermonde decomposition with nonnegative coefficients $\alpha_{j}\geq 0$, which means $\mathcal{B}$ is a complete Hankel tensor and hence CD. ∎
5.4 A Hankel tensor that is PSD but not CD
[…]
Let $m=4$ and $n=2$,
$\mathcal{A}=\sum_{j=1}^{3}\alpha_{j}\mathbf{u}_{j}^{4},$
where $\alpha_{1}=\alpha_{2}=1$, $\alpha_{3}=-0.0001$, $\mathbf{u}_{1}=(1,1)^{\top}$, $\mathbf{u}_{2}=(1,1000)^{\top}$ and $\mathbf{u}_{3}=(1,0.0001)^{\top}$. Then $\mathcal{A}$ is a Hankel tensor.
$\mathcal{A}$ is PD and hence PSD. The quartic form of $\mathcal{A}$ is:
$\mathcal{A}\mathbf{x}^{4}=(x_{1}+x_{2})^{4}+(x_{1}+1000x_{2})^{4}-0.0001(x_{1}+0.0001x_{2})^{4}$
The H-eigenvalues of $\mathcal{A}$ are all positive, with the smallest H-eigenvalue approximately 0.9956. Therefore, $\mathcal{A}$ is PD and hence PSD.
$\mathcal{A}$ is not a CD tensor. By definition, a CD tensor must be expressible as a sum of $m$th powers of vectors with nonnegative coefficients. The given decomposition contains a negative coefficient ($\alpha_{3}=-0.0001$), which violates this requirement. Moreover, moment analysis confirms that there exists no alternative decomposition of $\mathcal{A}$ as a sum of fourth powers with nonnegative coefficients.
This aligns with Theorem 2.1(iv), which states that for $m=4$ and $n=2$, $CD_{4,2}$ is a proper subset of $SOS_{4,2}=PSD_{4,2}$, i.e., $CD_{4,2}\subsetneq SOS_{4,2}$.
5.5 A Hankel tensor that is PSD but not SOS
[…]
By Hilbert [3], a PSD tensor in $S_{m,2}$ is also SOS. Thus, we may consider an example that $m=n=4$. Let
$\mathcal{A}=\sum_{j=1}^{5}\alpha_{j}\mathbf{u}_{j}^{4},$
where $\alpha_{1}=\alpha_{2}=\alpha_{3}=\alpha_{4}=1$, $\alpha_{5}=-0.0001$, $\mathbf{u}_{1}=(1,1,1,1)^{\top}$, $\mathbf{u}_{2}=(1,10,100,1000)^{\top}$, $\mathbf{u}_{3}=(1,20,400,8000)^{\top}$, $\mathbf{u}_{4}=(1,50,2500,125000)^{\top}$, and $\mathbf{u}_{5}=(1,0.0001,0.00000001,0.000000000001)^{\top}$. Then $\mathcal{A}$ is a Hankel tensor.
$\mathcal{A}$ is PD. The quartic form of $\mathcal{A}$ is:
$\mathcal{A}\mathbf{x}^{4}=\sum_{j=1}^{4}(\mathbf{u}_{j}\cdot\mathbf{x})^{4}-0.0001(\mathbf{u}_{5}\cdot\mathbf{x})^{4}.$
We can show that the positive terms dominate the negative term. Note that: For any $\mathbf{x}\neq\mathbf{0}$, we have:
$\sum_{j=1}^{4}(\mathbf{u}_{j}\cdot\mathbf{x})^{4}\geq 4\|\mathbf{x}\|^{4},$
and
$|\mathbf{u}_{5}\cdot\mathbf{x}|^{4}\leq\|\mathbf{x}\|^{4}.$
So
$\mathcal{A}\mathbf{x}^{4}\geq 4\|\mathbf{x}\|^{4}-0.0001\|\mathbf{x}\|^{4}=3.9999\|\mathbf{x}\|^{4}>0.$
Therefore, $\mathcal{A}$ is PD.
$\mathcal{A}$ is not a CD tensor. By definition, a CD tensor must be expressible as a sum of $m$th powers of vectors with nonnegative coefficients. The given decomposition contains a negative coefficient ($\alpha_{5}=-0.0001$), which violates this requirement. Moreover, there exists no alternative decomposition of $\mathcal{A}$ as a sum of fourth powers with nonnegative coefficients.
$\mathcal{A}$ is not SOS. Using the scientific computing software MATLAB, equipped with the YALMIP and SEDUMI packages, we verified that tensor A is not SOS, since the corresponding SOS optimization problem has no feasible solution. This example answers the open question from Subsection 5.7.10 of [8] in the affirmative: an even order PSD Hankel tensor may not be SOS.
This example illustrates the strict inclusions $CD_{4,4}\subsetneq SOS_{4,4}\subsetneq PSD_{4,4}$ mentioned in Theorem 2.1(iv).
@@ 6 SOS Tensors
Note on Notation and Normalization
In this section, we work with the unnormalized monomial basis. Specifically, for homogeneous polynomials of degree $k$, we use the basis $\{\mathbf{x}^{\beta}\}$ without multinomial coefficients. The inner product is defined as:
$\mathcal{A}\bullet\mathcal{B}=\sum_{|\alpha|=m}a_{\alpha}b_{\alpha}.$
This convention simplifies the moment matrix and polynomial representations while maintaining mathematical consistency throughout this section.
@@ 6.1 Answers to Fundamental Questions
    No. The SOS cone is not self-dual in general. For example, when $m=4$ and $n=2$, we have $SOS_{4,2}=PSD_{4,2}$ and $SOS_{4,2}^{*}=CD_{4,2}$.
@@ 7 Final Remarks
  • We also answered an open question raised in Subsection 5.7.10 of [8], by giving a PSD but not SOS Hankel tensor.
@@
  • Identifying the conditions for a PSD Hankel tensor to be SOS,
  • Investigating the relations of several SOS tensor subclasses, such as even-order symmetric weakly diagonally dominated tensors, even-order M-tensors and their absolute tensors, even-order symmetric B0-tensors, even-order H-tensors with nonnegative diagonal entries, with CD tensors.
Acknowledgment We are thankful to Yannan Chen who verified the example in Subsection 5.5. This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), the R&D project of Pazhou Lab (Huangpu) (Grant no. 2023K0603), the Fundamental Research Funds for the Central Universities (Grant No. YWF-22-T-204), Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002).